the difference of 3 times the larger consecutive integer and the smaller consecutive integer is 7
can somebody work this out and show me how you got it ?

Answer by stanbon(75887)

the difference of 3 times the larger consecutive integer and the smaller consecutive integer is 7
---
smaller: x
larger: x+1
------------------
Equation:
3(x+1) - x = 7
-----
3x+3-x = 7
2x + 3 = 7
2x = 4
x = 2
=============
Cheers,
Stan H.
=============